# Comprehensive Study of $k$-essence Model: Dynamical System Analysis and Observational Constraints from Latest Type Ia Supernova and BAO Observations

### v2

@@ I Introduction
Testing the $k$-essence model against observational data is crucial to determine how well it aligns with the observations. To perform this analysis, we numerically solve the field and Friedmann equations, thereby obtaining the Hubble parameter as a function of redshift. Setting initial conditions and priors for the model parameters involves conducting a dynamical system analysis. Stability techniques within dynamical systems are invaluable tools for examining complex models, as they facilitate the identification of physically viable solutions by analyzing the existence and nature of critical points, thus constraining the model parameters [41, 42, 43, 44, 45, 46, 47, 48]. These critical points correspond to the various phases of the universe. Some of the critical points are practically useful as they may have a connection to the late-time acceleration of the universe.
@@
In this study, we analyze the $k$-essence scalar field with the Lagrangian as $\mathcal{L}_{\phi}=V(\phi)(-X+X^{2})$, where $X=-\frac{1}{2}g^{\mu\nu}\nabla_{\mu}\phi\nabla_{\nu}\phi$, with two types of potentials: (i) inverse square potential $V(\phi)\propto\frac{1}{\phi^{2}}$ and (ii) exponential potential $V(\phi)\propto e^{\phi}$. We test these models using observational datasets including Cosmic Chronometer (CC), Baryonic Acoustic Oscillation (BAO), Dark Energy Survey Supernova 5YR (DES-SN), and SH0ES-calibrated Pantheon samples (Pantheon+SH0ES) of Type Ia supernovae. While extensive research has been conducted on the $k$-essence field with the inverse square potential, there is limited literature addressing the exponential potential within the Lagrangian framework specified above. Moreover, studies on pure kinetic $k$-essence fields are sparse [49, 50]. Our analysis indicates that the exponential potential model can exhibit a late-time cosmic accelerating phase, preceded by an attracting matter phase that persists for a brief period.
@@
The paper is organized as follows: In Sec. II, we analyze the $k$-essence model with inverse square and exponential type potentials from a dynamical system perspective and present the numerical evolution of the cosmological parameters. The data analysis of the $k$-essence models is discussed in Sec. III. Finally, we present our conclusions in Sec. IV.
II K-essence field
@@
$S=\int d^{4}x\sqrt{-g}\bigg{(}\frac{R}{2\kappa^{2}}-\mathcal{L}_{\phi}(X,\phi)+\mathcal{L}_{m}\bigg{)},$ (1)
@@
Here, the parameters $\delta$ and $\beta$ are dimensionless constants, whereas $H_{0}$ refers to a constant whose dimension is the same as the Hubble parameter. The other parameter $V_{0}$ is a dimensional constant with mass dimension four. We have considered the exponential potential because this type of potential is usually studied with quintessential types of scalar fields [23]. However, very few studies have been conducted with the $k$-essence field [52, 53, 54], particularly with the Lagrangian form $\mathcal{L}_{\phi}=-V(\phi)F(X)$. Hence, we will carry out a detailed study of the dynamics of the $k$-essence field with both potentials. To study the following scenario, we will take the particular form of the kinetic function as given by:
$F(X)=-X+X^{2}.$ (4)
@@
$T_{\mu\nu}^{(\phi)}=-{\mathcal{L}}_{,X}\,(\partial_{\mu}\phi)(\partial_{\nu}\phi)-g_{\mu\nu}\,{\mathcal{L}}\,.$ (5)
@@
$\rho_{\phi}=\mathcal{L}-2X\mathcal{L}_{,X}\quad\text{and}\quad P_{\phi}=-\mathcal{L}\,.$ (6)
@@
$\begin{split}\dot{\rho}_{m}+3H(\rho_{m}+P_{m})&=0,\\ \dot{\rho}_{\phi}+3H(\rho_{\phi}+P_{\phi})&=0.\end{split}$ (7)
@@
II.1 Stability of the system in the presence of Inverse square potential
@@
$x^{2}=\frac{\kappa^{2}V\dot{\phi}^{2}}{6H^{2}},\quad y^{2}=\frac{\kappa^{2}V\dot{\phi}^{4}}{4H^{2}},\quad\Omega_{m}=\frac{\kappa^{2}\rho_{m}}{3H^{2}},\quad\lambda=\frac{-V_{,\phi}}{\kappa V^{3/2}},\quad\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}.$ (11)
Although a large section of the literature has been published discussing the dynamics corresponding to the inverse square potential, we are explicitly showing it here for the sake of consistency. To analyze the dynamics, we choose a different dynamical setup to comprehend the dynamics of the phase space. While the dynamical variables are defined differently, the results we obtained match those of previous studies. We choose the variables corresponding to the field sector such that the Hubble equation takes a similar form to the quintessence field as presented in Ref.  [9]. The primary variables associated with the $k$-essence field are $(x,y>0)$, and are assumed to be positive. The variables $(\Omega_{\phi},\Omega_{m})$ are secondary variables as they can be expressed in terms of primary variables. The other variable, $\lambda$, is called the slope of the potential, which in this case depends on $\delta$ and becomes a constant. It should be noted that the dynamical variables defined here will not be valid for all types of potential. For instance, the dynamical variables chosen for the exponential potential in the next section in Eq. (23), although using similar notations, have definitions that are quite different from those used here. Following the Hubble equation Eq. (9), the primary variables are $(x,y,\Omega_{m})$ constrained as:
$0\leq\Omega_{m}=1-\Omega_{\phi}\leq 1,\quad 0\leq\Omega_{\phi}\leq 1$ (12)
@@
$\Omega_{\phi}=-x^{2}+y^{2},\quad\omega_{\rm eff}=\frac{P_{m}+P_{\phi}}{\rho_{m}+\rho_{\phi}}=\frac{y^{2}}{3}-x^{2}.$ (13)
@@
$\frac{\dot{H}}{H^{2}}=\frac{3}{2}\left(x^{2}-\frac{y^{2}}{3}-1\right).$ (14)
@@
$\begin{split}x^{\prime}&=-\frac{\dot{H}}{H^{2}}x+\frac{1}{2}\sqrt{6}(-\lambda)x^{2}+\frac{-\frac{1}{2}\sqrt{6}\lambda x^{2}-\frac{2y^{2}}{x}+3x+\frac{1}{2}\sqrt{6}\lambda y^{2}}{\frac{2y^{2}}{x^{2}}-1},\\ y^{\prime}&=-\frac{\dot{H}}{H^{2}}y+\frac{-\frac{4y^{3}}{x^{2}}+\frac{6\lambda y^{3}}{\sqrt{6}x}-\sqrt{6}\lambda xy+6y}{\frac{2y^{2}}{x^{2}}-1}+\frac{1}{2}\left(-\sqrt{6}\right)\lambda xy\ .\end{split}$ (15)
Here, the time derivative of the dynamical variables is closed and requires no extra variables, as a result, the phase space of the system becomes $2D$. The autonomous equations $(x^{\prime},y^{\prime})$ present the invariant submanifold $(x=0,y=0)$, hence the trajectories that originate in $x,y>0$ cannot cross the $x=y=0$ plane. For the inverse square type potential, the derivative of the potential yields:
$\lambda=2/\delta.$ (16)
@@
$J_{ij}=\frac{\partial x_{i}^{\prime}}{\partial x_{j}}\bigg{|}_{\bm{x}_{*}},$ (17)
@@
  • Point $P_{5}$: At this point, both field variables vanish, consequently, the field fractional density becomes zero. The fluid density dominates and the effective equation of state becomes 00, signifying matter-dominated type behavior. However, upon determining the Jacobian matrix, the corresponding eigenvalues yield negative values, making it stable.
@@
The phase space of the system has been plotted for $\lambda=0.2$ in Fig. [2], where the gray region shows the constraint on field energy density, $(0\leq\Omega_{\phi}\leq 1)$. The green region shows the accelerating characteristics, $(-1\leq\omega_{\rm eff}<-1/3)$, when field energy density dominates $0.5<\Omega_{\phi}\leq 1$, and the phantom region $(-1.5<\omega_{\rm eff}<-1)$ is shown in red color. From the phase space, it becomes clear that the trajectories originating from $P_{1}$ (where near $P_{1}$ we have radiation-type characteristics) get attracted towards both $P_{3}$ and $P_{4}$. Those trajectories whose slopes are positive, $x>0$ are attracted towards $P_{4}$, while the trajectories with negative slopes, $x<0$, get attracted towards $P_{3}$. Depending on the choice of $\lambda$, either $P_{3}$ or $P_{4}$ shows accelerating characteristics. There are a few trajectories which are not constrained by the Hubble equation and are not attracted towards $P_{3,4}$; they get attracted to $P_{5}$, which shows an attractor matter regime. Note that point $P_{2}$ cannot be considered as a physically viable point, since $\Omega_{m}$ becomes negative for $\lambda^{2}<3$. The trajectory that lies outside the constrained region shown in blue cannot be considered physically viable, as it violates the essential constraints imposed on the system by the Hubble equation, Eq. (12). In Fig. [3], we numerically evolved the system by varying the initial conditions and model parameters in the range of
$0.8\leq x_{0}\leq 1.12,\quad 0.1\leq\Omega_{m0}\leq 0.31,\quad 0.01\leq\lambda\leq 0.3,$ (18)
by generating random numbers. Here, the subscript zero denotes the initial conditions at $N=\tilde{z}=0$. We have also plotted the evolution of the Hubble parameter $(H)$ in km/s/Mpc and distance modulus, $\mu$, against redshift, $\tilde{z}=1/(1+a)$, where $\mu$ is defined as
$\mu=5\log_{10}(d_{L})+25,$ (19)
@@
$y_{0}=\sqrt{1-\Omega_{m0}+x_{0}^{2}}\ .$ (21)
In the figure, we can see that the initial conditions with $x_{0}>0$ produces values of $x$ that remain positive; hence, the system stabilizes at point $P_{4}$ and generates an accelerating solution. We see that although during the late-time epoch, the field energy density dominates, but in the past epoch, the curves corresponding to the Hubble evolution show deviation from the data points. However, there is no substantial difference that can be accounted for in the distance modulus plot. In the subsequent sec. III, we will conduct statistical analysis to obtain the best fit values of the dynamical variables and model parameters.
II.2 Stability of the system in the presence of exponential potential
@@
$x=\dot{\phi},\quad y=\phi H_{0},\quad z=\frac{H-H_{0}}{H+H_{0}}\,,$ (23)
@@
$-\infty<x<\infty,\quad-\infty<y<\infty,$ (24)
@@
$z=\begin{cases}-1&H\ll H_{0},\\ 0&H=H_{0},\\ 1&H\to+\infty\ .\end{cases}$ (25)
@@
$\Omega_{\phi}:=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}=\frac{\mathcal{M}_{0}}{3}\left(\frac{1-z}{1+z}\right)^{2}e^{\beta y}\left(\frac{3x^{4}}{4}-\frac{x^{2}}{2}\right),\quad\Omega_{m}=\frac{\kappa^{2}\rho_{m}}{3H^{2}}\,.$ (26)
@@
$\Omega_{m}=1-\Omega_{\phi},\quad 0\leq\Omega_{m}\leq 1,\quad 0\leq\Omega_{\phi}\leq 1.$ (27)
@@
$\omega_{\rm eff}=\frac{\mathcal{M}_{0}}{3}\left(\frac{1-z}{1+z}\right)^{2}e^{\beta y}\left(\frac{x^{4}}{4}-\frac{x^{2}}{2}\right).$ (28)
@@
$x^{\prime}$ $=$ $\frac{-3x^{3}+\frac{\beta\left(\frac{x^{2}}{2}-\frac{3x^{4}}{4}\right)(1-z)}{z+1}+3x}{3x^{2}-1},$ (29)
$y^{\prime}$ $=$ $\frac{x(1-z)}{z+1},$ (30)
$z^{\prime}$ $=$ $-\frac{3((1-z)(z+1))\left(\frac{1}{3}\mathcal{M}_{0}\left(\frac{x^{4}}{4}-\frac{x^{2}}{2}\right)\left(\frac{1-z}{z+1}\right)^{2}\exp(\beta y)+1\right)}{4}\,.$ (31)
@@
$0.21.01<x_{0}<1.25,\quad 0.3<\Omega_{m0}<0.31,\quad z_{0}=0,\quad 5<\mathcal{M}_{0}<9,\quad-0.2<\beta<$ 0.2 . (33)
Here, the subscript 00 indicates the value at the present epoch, for instance $x_{0}\equiv x(\tilde{N}=0)$. We have fixed $z_{0}=0$ due to its definition in Eq. (23). In this approach, we choose to vary $\Omega_{m0}$ instead of $y_{0}$ since, by following the Hubble constraint equation $\Omega_{m0}=1-\Omega_{\phi 0}$, one obtains the initial conditions for $y_{0}$. Note that this technique has also been adopted in the data analysis section III.
@@
III Data Analysis
Our subsequent objective involves performing data analysis to constrain the parameter space of the model. In order to achieve this goal, we need to derive an analytical expression for the Hubble parameter as a function of redshift. While this process is straightforward for the $\Lambda$CDM model, it poses challenges for $k$-essence models due to the inability to obtain an analytical solution to the Friedmann equation. Therefore, we opted for a dynamical systems approach, solving the coupled differential equations outlined in Eq. (15) to capture the evolution of the Hubble parameter. Here, we constraint the model parameters, utilizing the following data sets:

### v3

@@ 1 Introduction
Testing the $k$-essence model against observational data is crucial to determine how well it aligns with the observations. To perform this analysis, we numerically solve the field and Friedmann equations, thereby obtaining the Hubble parameter as a function of redshift. Setting initial conditions and priors for the model parameters involves conducting a dynamical system analysis. Stability techniques within dynamical systems are invaluable tools for examining complex models, as they facilitate the identification of physically viable solutions by analyzing the existence and nature of critical points, thus constraining the model parameters [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. These critical points correspond to the various phases of the universe. Some of the critical points are practically useful as they may have a connection to the late-time acceleration of the universe.
@@
In this study, we analyze the $k$-essence scalar field with the Lagrangian as $\mathcal{L}_{\phi}=V(\phi)(-X+X^{2})$, where $X=-\frac{1}{2}g^{\mu\nu}\nabla_{\mu}\phi\nabla_{\nu}\phi$, with two types of potentials: (i) inverse square potential $V(\phi)\propto\frac{1}{\phi^{2}}$ and (ii) exponential potential $V(\phi)\propto e^{C\phi}$ ($C$ is a dimensional constant). We test these models using observational datasets including Cosmic Chronometer (CC), Baryonic Acoustic Oscillation (BAO), Dark Energy Survey Supernova 5YR (DES-SN), and SH0ES-calibrated Pantheon samples (Pantheon+SH0ES) of Type Ia supernovae. While extensive research has been conducted on the $k$-essence field with the inverse square potential, there is limited literature addressing the exponential potential within the Lagrangian framework specified above. Moreover, studies on pure kinetic $k$-essence fields are sparse [61, 62]. Our analysis indicates that the exponential potential model can exhibit a late-time cosmic accelerating phase, preceded by a matter dominated phase that persists for a brief period.
@@
The paper is organized as follows: In Sec. 2, we analyze the $k$-essence model with inverse square and exponential type potentials from a dynamical system perspective and present the numerical evolution of the cosmological parameters. The data analysis of the $k$-essence models is discussed in Sec. 3. Finally, we present our conclusions in Sec. 4.
2 K-essence field
@@
$S=\int d^{4}x\sqrt{-g}\bigg{(}\frac{R}{2\kappa^{2}}-\mathcal{L}_{\phi}(X,\phi)+\mathcal{L}_{m}\bigg{)},$ (2.1)
@@
Here, the parameters $\delta$ and $\beta$ are dimensionless constants, whereas $H_{0}$ refers to a constant whose dimension is the same as the Hubble parameter. The other parameter $V_{0}$ is a dimensional constant with mass dimension four. We have considered the exponential potential because this type of potential is usually studied with quintessential types of scalar fields [23]. However very few studies, involving statistical analysis, have been conducted with the $k$-essence field [64, 65, 66], particularly with the Lagrangian form $\mathcal{L}_{\phi}=-V(\phi)F(X)$. Hence, we will carry out a detailed study of the dynamics of the $k$-essence field with both potentials. To study the following scenario, we will take the particular form of the kinetic function as given by:
$F(X)=-X+X^{2}.$ (2.4)
@@
$T_{\mu\nu}^{(\phi)}=-{\mathcal{L}}_{,X}\,(\partial_{\mu}\phi)(\partial_{\nu}\phi)-g_{\mu\nu}\,{\mathcal{L}}\,.$ (2.5)
@@
$\rho_{\phi}=\mathcal{L}-2X\mathcal{L}_{,X}\quad\text{and}\quad P_{\phi}=-\mathcal{L}\,.$ (2.6)
@@
$\begin{split}\dot{\rho}_{m}+3H(\rho_{m}+P_{m})&=0,\\ \dot{\rho}_{\phi}+3H(\rho_{\phi}+P_{\phi})&=0.\end{split}$ (2.7)
@@
2.1 Stability of the system in the presence of Inverse square potential
@@
$x^{2}=\frac{\kappa^{2}V\dot{\phi}^{2}}{6H^{2}},\quad y^{2}=\frac{\kappa^{2}V\dot{\phi}^{4}}{4H^{2}},\quad\Omega_{m}=\frac{\kappa^{2}\rho_{m}}{3H^{2}},\quad\lambda=\frac{-V_{,\phi}}{\kappa V^{3/2}},\quad\Omega_{\phi}=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}.$ (2.11)
Although a large section of the literature has been published discussing the dynamics corresponding to the inverse square potential, we are explicitly showing it here for the sake of consistency. To analyze the dynamics, we choose a different dynamical setup to comprehend the dynamics of the phase space. While the dynamical variables are defined differently, the results we obtained match those of previous studies. We choose the variables corresponding to the field sector such that the Hubble equation takes a similar form to the quintessence field as presented in Ref.  [9]. The primary variables associated with the $k$-essence field are $(x,y>0)$, and are assumed to be positive. The variables $(\Omega_{\phi},\Omega_{m})$ are secondary variables as they can be expressed in terms of primary variables. The other variable, $\lambda$, is called the slope of the potential, which in this case depends on $\delta$ and becomes a constant. It should be noted that the dynamical variables defined here will not be valid for all types of potential. For instance, the dynamical variables chosen for the exponential potential in the next section in Eq. (2.23), although using similar notations, have definitions that are quite different from those used here. Following the Hubble equation Eq. (2.9), the primary variables are $(x,y,\Omega_{m})$ constrained as:
$0\leq\Omega_{m}=1-\Omega_{\phi}\leq 1,\quad 0\leq\Omega_{\phi}\leq 1$ (2.12)
@@
$\Omega_{\phi}=-x^{2}+y^{2},\quad\omega_{\rm eff}=\frac{P_{m}+P_{\phi}}{\rho_{m}+\rho_{\phi}}=\frac{y^{2}}{3}-x^{2}.$ (2.13)
@@
$\frac{\dot{H}}{H^{2}}=\frac{3}{2}\left(x^{2}-\frac{y^{2}}{3}-1\right).$ (2.14)
@@
$\begin{split}x^{\prime}&=-\frac{\dot{H}}{H^{2}}x-\frac{1}{2}\sqrt{6}\lambda x^{2}+\frac{-\frac{1}{2}\sqrt{6}\lambda x^{2}-\frac{2y^{2}}{x}+3x+\frac{1}{2}\sqrt{6}\lambda y^{2}}{\frac{2y^{2}}{x^{2}}-1},\\ y^{\prime}&=-\frac{\dot{H}}{H^{2}}y+\frac{-\frac{4y^{3}}{x^{2}}+\frac{6\lambda y^{3}}{\sqrt{6}x}-\sqrt{6}\lambda xy+6y}{\frac{2y^{2}}{x^{2}}-1}-\frac{1}{2}\sqrt{6}\lambda xy\ .\end{split}$ (2.15)
Here, the time derivative of the dynamical variables is closed and requires no extra variables, as a result, the phase space of the system becomes $2D$. The autonomous equations $(x^{\prime},y^{\prime})$ present the invariant submanifold $(x=0,y=0)$, hence the trajectories that originate in $x,y>0$ cannot go out from this region. For the inverse square type potential, the derivative of the potential yields:
$\lambda=2/\delta.$ (2.16)
@@
$J_{ij}=\frac{\partial x_{i}^{\prime}}{\partial x_{j}}\bigg{|}_{\bm{x}_{*}},$ (2.17)
@@
  • Point $P_{5}$: At this point, both field variables vanish, consequently, the field fractional density becomes zero. The fluid energy density dominates and the effective equation of state becomes 00, signifying a state similar to a matter-dominated phase. However, upon determining the Jacobian matrix, the corresponding eigenvalues yield negative values, making it stable.
@@
The phase space of the system has been plotted for $\lambda=0.2$ in Fig. [2], where the gray region shows the constraint on field energy density, $(0\leq\Omega_{\phi}\leq 1)$. The green region shows the accelerating characteristics, $(-1\leq\omega_{\rm eff}<-1/3)$, when field energy density dominates $0.5<\Omega_{\phi}\leq 1$, and the phantom region $(-1.5<\omega_{\rm eff}<-1)$ is shown in red color. From the phase space, it becomes clear that the trajectories originating from $P_{1}$ (where near $P_{1}$ we have radiation-type characteristics) get attracted towards both $P_{3}$ and $P_{4}$. Those trajectories whose slopes are positive, $x>0$ are attracted towards $P_{4}$, while the trajectories with negative slopes, $x<0$, get attracted towards $P_{3}$. Depending on the choice of $\lambda$, either $P_{3}$ or $P_{4}$ shows accelerating characteristics. There are a few trajectories which are not constrained by the Hubble equation and are not attracted towards $P_{3,4}$; they get attracted to $P_{5}$, which shows an attractor matter regime. Note that point $P_{2}$ cannot be considered as a physically viable point, since $\Omega_{m}$ becomes negative for $\lambda^{2}<3$. The trajectory that lies outside the constrained region shown in blue cannot be considered physically viable, as it violates the essential constraints imposed on the system by the Hubble equation, Eq. (2.12). In Fig. [3], we numerically evolved the system by varying the initial conditions and model parameters in the range of
$0.8\leq x_{0}\leq 1.12,\quad 0.1\leq\Omega_{m0}\leq 0.31,\quad 0.01\leq\lambda\leq 0.3,$ (2.18)
by generating random numbers. Here, the subscript zero denotes the initial conditions at $N=\tilde{z}=0$. We have also plotted the evolution of the Hubble parameter $(H)$ in km/s/Mpc and distance modulus, $\mu$, against redshift, $a=1/(1+\tilde{z})$, where $\mu$ is defined as
$\mu=5\log_{10}(d_{L})+25,$ (2.19)
@@
$y_{0}=\sqrt{1-\Omega_{m0}+x_{0}^{2}}\ .$ (2.21)
In the figure, we can see that the initial conditions with $x_{0}>0$ produces values of $x$ that remain positive; hence, the system stabilizes at point $P_{4}$ and generates an accelerating solution. We see that although during the late-time epoch, the field energy density dominates, but in the past epoch, the curves corresponding to the Hubble evolution show deviation from the data points. However, there is no substantial difference that can be accounted for in the distance modulus plot. In addition to this analysis, it is essential to examine the critical points at infinity to determine if any viable fixed points exist. A comprehensive analysis of these points is provided in Appendix A. In the subsequent sec. 3, we will conduct statistical analysis to obtain the best fit values of the dynamical variables and model parameters.
2.2 Stability of the system in the presence of exponential potential
@@
$x=\dot{\phi},\quad y=\phi H_{0},\quad z=\frac{H-H_{0}}{H+H_{0}}\,,$ (2.23)
@@
$-\infty<x<\infty,\quad-\infty<y<\infty,$ (2.24)
@@
$z=\begin{cases}-1&H\ll H_{0},\\ 0&H=H_{0},\\ 1&H\to+\infty\ .\end{cases}$ (2.25)
@@
$\Omega_{\phi}:=\frac{\kappa^{2}\rho_{\phi}}{3H^{2}}=\frac{\mathcal{M}_{0}}{3}\left(\frac{1-z}{1+z}\right)^{2}e^{\beta y}\left(\frac{3x^{4}}{4}-\frac{x^{2}}{2}\right),\quad\Omega_{m}=\frac{\kappa^{2}\rho_{m}}{3H^{2}}\,.$ (2.26)
@@
$\Omega_{m}=1-\Omega_{\phi},\quad 0\leq\Omega_{m}\leq 1,\quad 0\leq\Omega_{\phi}\leq 1.$ (2.27)
@@
$\omega_{\rm eff}=\frac{\mathcal{M}_{0}}{3}\left(\frac{1-z}{1+z}\right)^{2}e^{\beta y}\left(\frac{x^{4}}{4}-\frac{x^{2}}{2}\right).$ (2.28)
@@
$x^{\prime}$ $=$ $\frac{-3x^{3}+\frac{\beta\left(\frac{x^{2}}{2}-\frac{3x^{4}}{4}\right)(1-z)}{z+1}+3x}{3x^{2}-1},$ (2.29)
$y^{\prime}$ $=$ $\frac{x(1-z)}{z+1},$ (2.30)
$z^{\prime}$ $=$ $-\frac{3((1-z)(z+1))\left(\frac{1}{3}\mathcal{M}_{0}\left(\frac{x^{4}}{4}-\frac{x^{2}}{2}\right)\left(\frac{1-z}{z+1}\right)^{2}\exp(\beta y)+1\right)}{4}\,.$ (2.31)
@@
Points $(x,y,z)$ $\Omega_{\phi}$ $\Omega_{m}$ $\omega_{\rm eff}$
$P_{1,2}$ $(0,\text{Any},\mp 1)$ 0 1 0
$P_{3,4}$ $(\mp 1,\text{Any},1)$ 0 1 0
Table 2: The nature of critical points corresponding to an exponential potential.
To analyze the system’s dynamics, we first determine the system’s critical points using $\tilde{N}$ and examine their stability. The critical points are presented in Table [2]. In the table - Any - shows that the $y$ coordinate can be chosen arbitrarily, the critical point analysis does not specify any particular values for the $y$-coordinate of the critical points. The system reveals four critical points where the matter fractional energy density dominates while the field fractional energy density vanishes. This results in an effective equation of state (EoS) that vanishes, indicating a matter-dominated regime.
Given the complexity of the system, it is unable to generate critical points that can directly constrain the model parameters $\beta$ and $\mathcal{M}_{0}$, as was possible in previous cases. Consequently, a purely numerical approach is required to evolve the system with respect to $\tilde{N}$ by varying initial conditions and model parameters to identify the best range of initial conditions that can replicate the different epochs of the universe.
Before moving on to the numerical evolution, we address a key concern for readers. The ranges for our dynamical variables are specified in Eq. (2.24). This range implies the potential existence of critical points at infinity, which could shed light on accelerating critical points. To address this, we utilize the Poincaré transformation for one of the variables, say $\bar{y}=y/\sqrt{1+y^{2}}$. Since $y$ can range from $\infty$- ∞ to +$\infty$+ ∞, this transformation maps these infinities to the interval $-1$ to $+1$, thereby compactifying the dynamical variable. With both $z$ and $\bar{y}$ confined within finite ranges, there is no need to compactify the variable $x$, as the Hubble constraint enforces valid physical solutions. Following this compactification approach, we can identify the critical points at infinity. In Appendix B, we demonstrate that the transformed variables do not exhibit any additional critical points beyond the current case, and, upon numerical simulation, the system dynamics remain consistent with our current scenario.
It is also worth noting that this compactification exercise can be avoided by a different line of reasoning. Although the dynamical variables can theoretically take any values, the Hubble constraint requires that the fractional energy densities (field or fluid) remain within $0\leq\Omega_{\phi}\leq 1$. From Eq. (2.26), we observe that the field energy density parameter depends on all variables, with $z$ already compactified. Therefore, the remaining values for $(x,y)$ must only yield finite densities. Any critical point that fails to satisfy this constraint cannot be considered physically viable.
@@
$0.21.01<x_{0}<1.25,\quad 0.3<\Omega_{m0}<0.31,\quad z_{0}=0,\quad 5<\mathcal{M}_{0}<9,\quad-0.2<\beta<$ 0.2 . (2.33)
Here, the subscript 00 indicates the value at the present epoch, for instance $x_{0}\equiv x(\tilde{N}=0)$. We have fixed $z_{0}=0$ due to its definition in Eq. (2.23). In this approach, we choose to vary $\Omega_{m0}$ instead of $y_{0}$ since, by following the Hubble constraint equation $\Omega_{m0}=1-\Omega_{\phi 0}$, one obtains the initial conditions for $y_{0}$. Note that this technique has also been adopted in the data analysis section 3.
@@
3 Data Analysis
Our subsequent objective involves performing data analysis to constrain the parameter space of the model. In order to achieve this goal, we need to derive an analytical expression for the Hubble parameter as a function of redshift. While this process is straightforward for the $\Lambda$CDM model, it poses challenges for $k$-essence models due to the inability to obtain an analytical solution to the Friedmann equation. Therefore, we opted for a dynamical systems approach, solving the coupled differential equations outlined in Eq. (2.15) to capture the evolution of the Hubble parameter. Here, we constraint the model parameters, utilizing the following data sets: